The velocity ,vms-1,of a fairground ride t seconds after it begins is given by v = (3e-t + 7)sin t

What is the acceleration of the ride 2 seconds after it begins?


1
Expert's answer
2021-06-07T15:51:04-0400
"v(t)=(3e^{-t}+7)\sin t"

"a(t)=v'(t)=-3e^{-t}\sin t+(3e^{-t}+7)\cos t"


"a(2)=-3e^{-2}\sin (2)+(3e^{-2}+7)\cos (2) (m\cdot s^{-2})"

"\approx-3.451166\ (m\cdot s^{-2})"